OpenStudy (anonymous):

how much of a (n)80% orange juice drink must be mixed with 3 gallons of a 30% orange juice mix to obtain a mixture that is 50% orange juice

OpenStudy (anonymous):

anybody i really need help answering this question @Data_LG2

OpenStudy (anonymous):

we need a variable

OpenStudy (anonymous):

say \(x\) is the amount of 80% orange juice, which will have \(.8x\) orange juice then add it to 3 gallons of 30% juice which will have \(30\times .3=9\) gallons of juice

OpenStudy (anonymous):

the total will be \(x+50\) gallons which you want to be 50% juice set \[.8x+9=.5(x+50)\] and solve for \(x\)

OpenStudy (anonymous):

x=53.3

OpenStudy (anonymous):

yah but i made a mistake

OpenStudy (anonymous):

should be\[.8x+9=.5(x+3)\]

OpenStudy (anonymous):

x=-25

OpenStudy (anonymous):

man i am all messed up hold on

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

not 30 gallons, 3 gallons!!

OpenStudy (anonymous):

haha yes 3 !

OpenStudy (anonymous):

\[.8x+.9=.5(x+3)\]

OpenStudy (anonymous):

so \(x=2\)

OpenStudy (anonymous):

okay so the answer to the question is 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you ! i

OpenStudy (anonymous):

i have another question if you could help me